sorry i made a typing error before
An infinite number of ordered pairs, including (0,1) (1,0) and (2,-1) satisfies both equations. Explain why.
x+y=1
2x+2y=2

Found 2 solutions by Alan3354, Mathtut:
Answer by Alan3354(69443)

sorry i made a typing error before
An infinite number of ordered pairs, including (0,1) (1,0) and (2,-1) satisfies both equations. Explain why.
x+y=1
2x+2y=2
----------------
The 2nd is eqn 1 multiplied by 2. If you graph it, you'll get the same line.
It's only one equation.

Answer by Mathtut(3670)

lol......yes , I pretty much figured that
:
x+y=1.........eq 1
2x+2y=2.......eq 2
:
I describe to you already what it takes to have infinite solutions. one equation has to be a multiple of the other
:
take eq 1 and multiply it by 2--->2(x+y-1)=2x+2y=2....this is the same as eq 2
:
this means they are the same line...basically one laying on top of the other.....that is why they have infinite solutions because every ordered pair that works in one works in the other
:
God Bless...Mathtut